OpenStudy (anonymous):

how would i rewrite 6 & 5/8 - 4 & 1/2 as an addition sign ?

OpenStudy (saifoo.khan):

\[\Large 4\frac{5}{8} - 4\frac{1}{2}\] Like that?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

its a siz instead of a 4 & i gotta make the sign into a positive..

OpenStudy (anonymous):

\[2 \frac{5}{8} \text{ + } \frac{1}{2} \]